OpenStudy (anonymous):

ok so I have no idea how to do this question.can anyone walk me through it? I tried it on wolframalpha just to see what it says the graph looks like but I have no idea how to on paper.Q: You are having a conference call with the CEO of a paper company. You have interpreted the number of trees cut down versus profit as the function P(x) = –x4 + x3 + 7x2 − x − 6. Describe to the CEO what the graph looks like and, in general, how to sketch the graph without using technology. Use complete sentences, and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0)

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

@m&msdoodle ?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@Agl202

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (anonymous):

@JesseReyna

OpenStudy (trojanpoem):

P(x) = Profits x number of trees cut at the minima of the function your company break down. the highest profit at the local maxima You could graph the function by finding the turning points, (2nd derivative) , local maxima and minima ( 1st derivative) and where the function is increasing and decreasing (1st derivative too) You will have an image of how the graph looks like other method would be by taking some points. like -4,3, -2,-1 , 0 , 1, 2, 3, 4 until you get full description of the graph.

OpenStudy (mathmale):

Note that your function, P(x) = –x^4 + x^3 + 7x^2 − x − 6, is of the fourth order. You could learn something significant about fourth order functions if you'd please graph y=x^4 first, and then y=-x^4. Which graph opens up? Which graph opens down? Which of these sample graphs is more likely to model the given function correctly?

OpenStudy (mathmale):

Can you finish this sentence? "as x increases, or as we move to the right, the graph of P(x) = –x^4 + x^3 + 7x^2 − x − 6 begins in the ____ quadrant, goes thru its vertical intercept ( ? , ? ), and finally ends in quadrant _____ "

OpenStudy (anonymous):

"as x increases, or as we move to the right, the graph of P(x) = –x^4 + x^3 + 7x^2 − x − 6 begins in the 3rd quadrant, goes thru its vertical intercept ( 0 , -6 ), and finally ends in quadrant 4 " @mathmale